# seg2med: a segmentation-based medical image generation framework using denoising diffusion probabilistic models

### v1

@@ II-F Evaluation Metrics
To determine the statistical significance of the improvements in the SSIM, MAE, and PSNR, we performed statistical hypothesis tests. We compared the DDPM with each other model and applied a Mann-Whitney U test as the metrics are non-normally distributed. Our null hypothesis states that the DDPM did not lead to improvements in either metric. The null hypothesis is rejected if p <0.05.
@@
III Results
III-A Experiment I: Generation of CT Images
Table II(d) provides the quantitative results for CT image generation achieved by each implemented model. The models were trained using the training datasets from the SynthRAD Pelvis and the Internal Abdomen, while the M2OLIE Abdomen consisted of ”unseen” patients for evaluation. The results indicate that, although DDPM shows no significant advantage over other models in validation on the SynthRAD Pelvis and the Internal Abdomen, it demonstrates a clear improvement for the M2OLIE Abdomen. Specifically, DDPM achieves the highest SSIM of $0.93\pm 0.02$, the lowest MAE of $40.05\pm 13.15$, a relatively high PSNR of $26.16\pm 1.62$, and the lowest standard deviation across all metrics, indicating its superior and more stable performance compared to other models. Overall, DDPM achieves the highest SSIM of $0.94\pm 0.02$, the lowest MAE of $36.79\pm 24.34$, the lowest FID of 3.62, although the highest PSNR is observed with UNet ($28.05\pm 4.12$).
Table III presents the results of the significance tests, demonstrating that DDPM significantly outperformed other models in SSIM and MAE metrics (p <0.05) but performed significantly worse than other models in terms of PSNR (p <0.05).
@@ III-A Experiment I: Generation of CT Images
To provide a comprehensive comparison of the four models, Figure 4 presents line charts and box plots of the metrics for synthetic CT images of test patients across all three datasets. In line charts, the x-axis represents the patient index, and the y-axis indicates the mean metric value for the generated CT images of each patient. The three datasets are distinguished by background colors. In the box plots, the x-axis represents the four models, and the y-axis shows the metric values. Additionally, the exact values for all evaluated patients are displayed as points, with each dataset distinguished by a unique point color. The line charts and box plots show that DDPM consistently delivers stable and excellent performance across all datasets, while the other models exhibit poorer performance for some patients, particularly in the new dataset (e.g., patients 41, 42, and 43). This observation is further supported by DDPM’s lowest standard deviation across all metrics in Table II(d).
@@
mean±$\pm$±std UNet CycleGAN pix2pix DDPM
@@
mean±$\pm$±std UNet CycleGAN pix2pix DDPM
@@
III-B Experiment II: Generation of MR Images
@@
III-C Experiment III: Conversion from Digital Phantoms to CT Images
@@
III-D Experiment IV: CT-MR Image Conversion
@@
III-E Experiment V: Segmentation Evaluation of Synthetic CT Images
@@ IV Discussion
IV-A Comparison to Published Medical Image Synthesis Methods
The proposed framework outperformed established generative models such as UNet, CycleGAN, and pix2pix in synthesizing pelvic and abdominal CT images (Figure 2). Specifically, the 2-channel DDPM maintained superior structural integrity and realistic presentation of organs such as the liver, spleen, and aorta as shown in the FOVs. The quantitative results in Table II(d) show that the DDPM achieved the highest SSIM of 0.94$\pm$ 0.02 and the lowest MAE of 36.79$\pm$24.34, demonstrating its alignment with real images. However, its relatively lower PSNR of 27.58$\pm$3.48 suggests room for further improvement in the denoising process.
@@ IV-A Comparison to Published Medical Image Synthesis Methods
IV-B Limitations
Despite its strengths, this framework has limitations. The reliance on 2D image generation is a limitation, as it results in non-smooth cross-sectional reconstructions in lateral and coronal planes, reducing the continuity between slices. Additionally, the sampling process is relatively slow, requiring approximately 30 seconds per 512×512 slice with 500 timesteps on an NVIDIA RTX A6000 GPU. Attempts to implement 3D-DDPM were constrained by GPU memory, allowing generation only for 128×128×128 volumes. Future work should focus on optimizing 3D generation to achieve a balance between inference speed and image quality. Furthermore, the framework lacks conditional generation capabilities, limiting its ability to synthesize images with pathological features. For instance, although the internal abdomen CT dataset includes 78 aortic dissection patients, identifying pathological features necessitates continuous 3D volumes, which was not supported by the current framework. Future efforts will aim to train the 3D-seg2med network with disease-specific conditions and evaluate synthetic pathological features through collaborations with clinicians and diagnostic AI tools. Lastly, the quantitative results for CT images converted from MR images were relatively poor, with a mean SSIM of 0.77$\pm$0.04, a PSNR of 15.53$\pm$1.47, and an MAE of 16.32$\pm$4.66. This performance discrepancy can be attributed to the lower generalizability of the DDPM-MR model, due to its training on a smaller dataset compared to the DDPM-CT model. To address this limitation, future work will focus on incorporating a larger and more diverse MR dataset to enhance the model’s robustness and performance. Additionally, further optimization of the training process and exploration of advanced techniques for handling limited data scenarios will be prioritized.
V Conclusion
@@
In this study, we introduced seg2med, an innovative image generation framework that combines segmentation, image processing, image synthesis using Denoising Diffusion Probabilistic Models (DDPM), and evaluation steps. Our framework effectively generates high-quality synthetic CT and MR images from segmentation masks derived from real patient data and XCAT digital phantoms. It achieves SSIM values exceeding 0.90 for CT images and 0.80 for MR images, along with FSIM values above 0.70 for digital phantom CT images. Additionally, with an overall FID score below 4.00 for CT image generation, the framework demonstrates strong generalizability and fidelity across diverse datasets. Furthermore, it successfully generated paired multimodal CT and MR images with consistent anatomical structures, highlighting its potential for multimodal imaging applications.
@@
Overall, this study highlights the potential of seg2med as a robust and versatile tool for medical image synthesis, offering promising applications in clinical imaging, data augmentation, and the development of advanced diagnostic and therapeutic algorithms. By addressing its current limitations, seg2med could play a pivotal role in shaping the future of medical imaging and AI-driven healthcare solutions.

### v2

@@ III-F Evaluation Metrics
To determine the statistical significance of the improvements in the SSIM, MAE, and PSNR, we performed statistical hypothesis tests. We compared the DDPM-CT with each other model and applied a Mann-Whitney U test as the metrics are non-normally distributed. Our null hypothesis states that the DDPM did not lead to improvements in either metric. The null hypothesis is rejected if p <0.05.
@@
IV Results
IV-A Experiment I: Generation of CT Images
Table II(d) provides the quantitative results for CT image generation achieved by each implemented model. The models were trained using the training datasets from the SynthRAD Pelvis and the Internal Abdomen, while the M2OLIE Abdomen consisted of ”unseen” patients for evaluation. The results indicate that, although DDPM-CT shows no significant advantage over other models in validation on the SynthRAD Pelvis and the Internal Abdomen, it demonstrates a clear improvement for the M2OLIE Abdomen. Specifically, DDPM-CT achieves the highest SSIM of $0.93\pm 0.02$, the lowest MAE of $40.05\pm 13.15$, a relatively high PSNR of $26.16\pm 1.62$, and the lowest standard deviation across all metrics, indicating its superior and more stable performance compared to other models. Overall, DDPM-CT achieves the highest SSIM of $0.94\pm 0.02$, the lowest MAE of $36.79\pm 24.34$, the lowest FID of 3.62, although the highest PSNR is observed with UNet ($28.05\pm 4.12$).
Table III presents the results of the significance tests, demonstrating that DDPM-CT significantly outperformed other models in SSIM and MAE metrics (p <0.05) but performed significantly worse than other models in terms of PSNR (p <0.05).
@@ IV-A Experiment I: Generation of CT Images
To provide a comprehensive comparison of the four models, Figure 4 presents line charts and box plots of the metrics for synthetic CT images of test patients across all three datasets. In line charts, the x-axis represents the patient index, and the y-axis indicates the mean metric value for the generated CT images of each patient. The three datasets are distinguished by background colors. In the box plots, the x-axis represents the four models, and the y-axis shows the metric values. Additionally, the exact values for all evaluated patients are displayed as points, with each dataset distinguished by a unique point color. The line charts and box plots show that DDPM-CT consistently delivers stable and excellent performance across all datasets, while the other models exhibit poorer performance for some patients, particularly in the new dataset (e.g., patients 41, 42, and 43). This observation is further supported by DDPM-CT’s lowest standard deviation across all metrics in Table II(d).
@@
mean±$\pm$±std UNet CycleGAN pix2pix DDPM-CT
@@
mean±$\pm$±std UNet CycleGAN pix2pix DDPM-CT
@@
IV-B Experiment II: Generation of MR Images
@@
IV-C Experiment III: Conversion from Digital Phantoms to CT Images
@@
IV-D Experiment IV: CT-MR Image Conversion
@@
IV-E Experiment V: Segmentation Evaluation of Synthetic CT Images
@@ V Discussion
V-A Comparison to Published Medical Image Synthesis Methods
The proposed framework outperformed established generative models such as UNet, CycleGAN, and pix2pix in synthesizing pelvic and abdominal CT images (Figure 2). Specifically, the 2-channel DDPM maintained superior structural integrity and realistic presentation of organs such as the liver, spleen, and aorta as shown in the FOVs. The quantitative results in Table II(d) show that the DDPM-CT achieved the highest SSIM of 0.94$\pm$ 0.02 and the lowest MAE of 36.79$\pm$24.34, demonstrating its alignment with real images. However, its relatively lower PSNR of 27.58$\pm$3.48 suggests room for further improvement in the denoising process.
@@ V-A Comparison to Published Medical Image Synthesis Methods
V-B Limitations and Future Work
Despite its strengths, the proposed framework still has several limitations.
First, the reliance on 2D image generation leads to non-smooth reconstructions across sagittal and coronal planes, reducing inter-slice continuity. Although we attempted to extend the model to 3D-DDPM, GPU memory constraints restricted the training to small volumes (128×128×128). Future work will focus on optimizing efficient 3D architectures to balance quality and speed.
Second, the performance of MR image synthesis lags behind that of CT, particularly in sequences such as T1 VIBE IN and T1 VIBE OPP. This is attributed to two factors: (1) the MR training dataset is smaller in size compared to CT, and (2) MR data exhibits greater heterogeneity, involving multiple acquisition sequences with distinct signal characteristics. Our experiments revealed modality confusion effects, where the model struggled to distinguish between similar MR sequences. In contrast, synthesis performance was notably better on the more consistent and abundant T1 GRE sequence from SynthRAD. To mitigate this, we plan to incorporate sequence-specific embedding strategies and targeted data augmentation for rare MR sequences.
Third, in this study, separate DDPMs were trained for CT and MR synthesis. However, our recent experiments suggest that a unified model, conditioned on modality labels via class embeddings injected into the time embedding, can achieve comparable or even superior performance. This is based on the premise that both CT and MR share a common anatomical foundation. Unifying the models could enhance anatomical consistency and generalization, especially when trained on multi-modal datasets. We are actively exploring this direction and plan to present full results in future work.
Fourth, effective modality separation becomes critical in unified models. Our current models lack explicit mechanisms to disentangle modality-specific characteristics. As such, we are investigating contrastive learning approaches, such as the NT-Xent loss [29], to regularize the model based on the hypothesis that the denoising function $\epsilon_{\theta}$ exhibits modality-dependent structure. Preliminary results indicate this strategy could improve both interpretability and generation fidelity for underrepresented modalities like MR.
Lastly, the current physical prior simulation in PhysioSynth depends on organ-wise lookup tables and pre-defined MR signal equations. Although effective, this hand-crafted process may lack generalizability for new modalities or patient cohorts. In future work, we aim to develop a lightweight neural network to estimate modality-specific priors from segmentation masks in a learnable, data-driven manner.
VI Conclusion
@@
In this study, we introduced seg2med, an innovative image generation framework that combines segmentation, image processing, image synthesis using DDPM, and evaluation steps. Our framework effectively generates high-quality synthetic CT and MR images from segmentation masks derived from real patient data and XCAT digital phantoms. It achieves SSIM values exceeding 0.90 for CT images and 0.80 for MR images, along with FSIM values above 0.70 for digital phantom CT images. Additionally, with an overall FID score below 4.00 for CT image generation, the framework demonstrates strong generalizability and fidelity across diverse datasets. Furthermore, it successfully generated paired multimodal CT and MR images with consistent anatomical structures, highlighting its potential for multimodal imaging applications.
@@
Overall, this study highlights the potential of seg2med as a robust and versatile tool for medical image synthesis, offering promising applications in clinical imaging, data augmentation, and the development of advanced diagnostic and therapeutic algorithms. With further improvements, seg2med could unlock new possibilities in cross-modality image synthesis and fusion, making it a valuable tool for comprehensive AI-driven medical diagnosis.